Solve for x and y in the following set of simultaneous differential equations by using D-operator methods: (D-2)x + Dy = 10sin2t

Dx + (D+2)y = 0


1
Expert's answer
2021-07-26T07:55:52-0400

Let's write the given system of ODE in the matrix form

"\begin{pmatrix}
D-2 & D\\
D & D+2
\end{pmatrix}
\begin{pmatrix} x\\ y \end{pmatrix} = \begin{pmatrix} 10 \sin 2t\\ 0 \end{pmatrix}"

Where "D" is the differential operator "D=\frac{d}{dt}".

The determinant of this system is "(D-2)(D+2)=D^2-4" is invertible, therefore the matrix of the system is invertible too.


For arbitrary invertible matrix 2x2 we have

"\begin{pmatrix}
a & b\\
c & d
\end{pmatrix}^{-1} = \frac{1}{ad-bc}\begin{pmatrix}
d & -c\\
-b & a
\end{pmatrix}"

Therefore,

"\begin{pmatrix}
D-2 & D\\
D & D+2
\end{pmatrix}^{-1}=-\frac{1}{4}\begin{pmatrix}
D+2 & -D\\
-D & D-2
\end{pmatrix}"

and

"\begin{pmatrix} x\\ y \end{pmatrix} =\begin{pmatrix}
D-2 & D\\
D & D+2
\end{pmatrix}^{-1}\begin{pmatrix} 10 \sin 2t\\ 0 \end{pmatrix} =-\frac{1}{4}\begin{pmatrix}
D+2 & -D\\
-D & D-2
\end{pmatrix}\begin{pmatrix} 10 \sin 2t\\ 0 \end{pmatrix}"


"x=-\frac{1}{4}(D+2)10 \sin 2t=-\frac{1}{4}(20\cos2t+20\sin 2t)=-5\cos2t-5\sin 2t"

"y=-\frac{1}{4}D(10\sin 2t)=-5 \cos 2t"


Answer. "x=-5\cos2t-5\sin 2t", "y=-5 \cos 2t".